Find the sum of n terms of the series 1/(√1+√3)+ 1/(√3+√5)+1/(√5+√7)+........upto n terms

Grade:11

2 Answers

Mohit Dhaka
7 years ago
Answer is nChange every term in difference of 2 terms by multiplying and dividing by 2After this- 2 in numeratorcan be written as=3-1=5-3=7-5....=2n+1-2n-1 then rationalise every term . You will fond that starting term get cancelled by ending term of succedding groupof term=1/2(2n+1-1)=n Here i cannot write the answer as i do not know how to write squre root
Hitesh
5 years ago
N th term is 1 / √2n-1+√2n+1.
On rationalizing we get
√2n+1-√2n+-1 / 2
We can write it as √2n+1/ 2 - √2n-1 / 2
Star puting values in the above eq. 
Tema will start canceling out and we get 
√2n+1 / 2 - 1/2 
= 1/2 {(√2n+1)-1}